# Outlyingness Scores with Cluster Catch Digraphs

### v1

@@ 3 Monte Carlo Experiments Under Random Cluster Process
In our previous paper [15], we have simulated three Neyman-Scott cluster processes (the Matérn, Thomas, and mixed cluster processes) to introduce randomness in both the location and size of clusters, The four CCD-based outlier detection methods were compared to several established state-of-the-art approaches, including Local Outlier Factor (LOF) [1], Density Based Spatial Clustering of Applications with Noise (DBSCAN) [2], the Minimal Spanning Tree (MST) Method [18], Outlier Detection using In-degree Number (ODIN) [4], and isolation Forest (iForest) [9]. Simulation analysis shows LOF achieved the highest $F_{2}$ scores in over half of the simulation settings, and the SUN-MCCD method demonstrated the second-best overall performance while simultaneously providing clustering results. Moreover, it is worthy noting that the SUN-MCCD method outperforms other cluster-based outlier detection methods by substantial margins, which makes it a decent choice under most simulation settings.
@@
We assess the performance of the four outlying scores by repeating the random cluster process in Section 6 of [15]. For comparison, This analysis included the four CCD-based methods and existing outlier detection methods for comparison. To Expand on the previous evaluation, we increase the maximum dimensions from 20 to 100 to evaluate all algorithmic performance in high-dimensional space. As noted in Section 2, IOS can be affected by the masking problem. This issue is exacerbated as the data size $n$ goes from 50 to 1000 in low-dimensional settings due to the increase in outlier intensity. To address this, we introduce a parameter, $S_{min}$, set to 0.04, for the two IOSs (excluding OOSs since they are not cluster-based), filtering the clusters with sizes below $S_{min}$. We use the same thresholds in Tables 1 and 2 for the Matérn cluster process and the Thomas cluster process, respectively. For the mixed cluster process, we set the thresholds as the mean of the values from the two tables. For example, when $d=100$, the thresholds of RKCCD-OOS are 13 (Table 1) and 10 (Table 2), resulting in a threshold of 11.5 for the mixed cluster process.
@@
For LOF, the lower and upper bound of $k$ are 11 and 30, respectively. The highest LOF is identified for each point, with a threshold to 1.5. In DBSCAN, the outlier percentage is assumed known to get an appropriate cutoff value for the $4$-dist, serving as a threshold for outliers. For MST, inconsistent edges are identified using a threshold of 1.2, and clusters smaller than 2% of the data set are labeled as outliers. As for ODIN, we set the input parameters $k$ and $T$ to $n^{0.5}$ and $n^{0.33}$, where $n$ is the data size; iForests are constructed with 1000 iTrees, each with a sub-sample size of 256, a threshold of 0.55 is applied to capture the outliers.
@@
RKCCD-OOS 0.907 0.913 0.897 0.932 0.830 0.927 0.754 0.986 0.543 0.985 0.314 0.907 0.311 0.905
RKCCD-IOS 0.955 0.972 0.952 0.977 0.976 0.966 0.991 0.984 0.997 0.990 0.997 1.000 0.999 1.000
UNCCD-OOS 0.700 0.970 0.824 0.983 0.853 0.990 0.767 0.981 0.696 0.983 0.424 0.928 0.314 0.904
UNCCD-IOS 0.949 0.953 0.955 0.969 0.958 0.979 0.985 0.978 0.988 0.983 0.998 0.999 1.000 1.000
@@
Table 7: The TPRs and TNRs of selected outlier detection methods and OSs under a Matérn cluster process (the simulation setting I in Section 6 of [15]).
@@
RKCCD-OOS 0.910 0.657 0.915 0.774 0.879 0.735 0.870 0.764 0.764 0.568 0.611 0.262 0.608 0.259
RKCCD-IOS 0.964 0.838 0.965 0.898 0.971 0.909 0.988 0.959 0.994 0.974 0.999 0.997 1.000 0.999
UNCCD-OOS 0.835 0.627 0.904 0.811 0.922 0.855 0.874 0.768 0.840 0.702 0.676 0.402 0.609 0.261
UNCCD-IOS 0.951 0.770 0.962 0.884 0.969 0.916 0.982 0.938 0.986 0.951 0.999 0.997 1.000 1.000
@@
Table 8: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under a Matérn cluster process (the simulation setting I in Section 6 of [15]).
@@
RKCCD-OOS 0.758 0.917 0.727 0.937 0.800 0.928 0.861 0.926 0.724 0.909 0.540 0.921 0.547 0.943
RKCCD-IOS 0.891 0.951 0.918 0.936 0.941 0.919 0.943 0.897 0.808 0.997 0.912 0.878 0.924 0.820
UNCCD-OOS 0.546 0.946 0.661 0.954 0.851 0.942 0.830 0.920 0.798 0.930 0.573 0.890 0.548 0.943
UNCCD-IOS 0.867 0.953 0.897 0.949 0.948 0.925 0.899 0.963 0.917 0.916 0.927 0.938 0.924 0.823
@@
Table 9: The TPRs and TNRs of selected outlier detection methods and OSs under a Thomas cluster process (the simulation setting II in Section 6 of [15]).
@@
RKCCD-OOS 0.838 0.506 0.832 0.639 0.864 0.711 0.894 0.761 0.817 0.637 0.731 0.506 0.745 0.563
RKCCD-IOS 0.921 0.677 0.927 0.782 0.930 0.804 0.920 0.783 0.903 0.831 0.895 0.751 0.872 0.710
UNCCD-OOS 0.746 0.416 0.808 0.613 0.897 0.768 0.875 0.732 0.864 0.722 0.732 0.505 0.746 0.538
UNCCD-IOS 0.910 0.667 0.923 0.789 0.937 0.817 0.931 0.844 0.917 0.787 0.933 0.851 0.874 0.712
@@
Table 10: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under Thomas cluster process (the simulation setting II in Section 6 of [15]).
@@
RKCCD-OOS 0.755 0.943 0.756 0.959 0.749 0.974 0.817 0.964 0.636 0.954 0.473 0.792 0.367 0.808
RKCCD-IOS 0.928 0.950 0.922 0.959 0.886 0.949 0.868 0.937 0.969 0.880 0.997 0.950 1.000 0.994
UNCCD-OOS 0.700 0.970 0.824 0.983 0.853 0.990 0.766 0.981 0.696 0.983 0.424 0.928 0.314 0.904
@@
Table 11: The TPRs and TNRs of selected outlier detection methods and OSs under a mixed cluster process (the simulation setting III in Section 6 of [15]).
@@
RKCCD-OOS 0.849 0.583 0.858 0.735 0.862 0.735 0.891 0.780 0.795 0.612 0.633 0.326 0.588 0.242
RKCCD-IOS 0.939 0.729 0.941 0.833 0.918 0.804 0.903 0.767 0.925 0.770 0.974 0.913 0.997 0.989
UNCCD-OOS 0.835 0.627 0.904 0.811 0.922 0.855 0.874 0.768 0.840 0.708 0.676 0.402 0.609 0.261
@@
Table 12: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under a mixed cluster process (the simulation setting III in Section 6 of [15]).
@@
U-MCCDs 8 8 9 12 11 8 8 9 9 11 12 13 7 7 9 10 10 12 13 8 7
SU-MCCDs 4 5 5 6 10 7 8 3 3 7 11 12 7 9 4 5 6 11 11 8 7
UN-MCCDs 9 7 7 5 8 11 10 7 6 8 8 11 13 13 8 9 7 5 9 10 9
@@
RKCCD-OOS 10 11 12 11 12 13 13 11 10 10 5 9 10 8 11 11 11 7 10 13 13
RKCCD-IOS 2 2 2 1 1 1 2 5 5 4 4 1 4 4 6 3 8 9 5 2 2
UNCCD-OOS 11 9 8 10 9 12 12 12 12 6 9 6 11 11 10 6 4 8 7 11 11
UNCCD-IOS 7 3 1 2 2 1 1 6 4 3 2 3 1 3 3 2 1 1 1 1 1
LOF 1 1 6 8 5 6 5 1 1 1 1 2 3 2 1 1 2 3 4 4 4
DBSCAN 3 10 10 9 7 5 6 2 8 9 7 5 5 6 2 8 9 10 6 6 6
@@
ODIN 6 4 4 4 4 4 4 8 7 5 6 7 6 5 7 7 5 4 3 5 5
iForest 12 12 11 7 3 3 3 10 11 12 10 4 2 1 12 12 12 6 2 3 3
Table 13: The rankings (by $F_{2}$-scores) of all the methods under each simulation setting of this section, top 3 are highlighted in bold.
Considering the two OOSs, while achieving high TNRs, both deliver lower TPRs and $F_{2}$-scores, ranking around $10^{th}$ place out of 13 methods under most simulation settings. For instance, the TPRs of RKCCD-OOS are 0.755, 0.756, 0.749, 0.817, 0.636, 0.473, and 0.367 under the Matérn cluster process, substantially lower than most other methods. The reason stems from the limitation of OOS: OOS measures the outlyingness of a point by comparing its vicinity density to its outbound neighbors. This approach is not robust to the masking problem. Moreover, the masking problem is more severe under the simulations with the three Neyman-Scott cluster processes compared to those with distinct clusters, as the random cluster processes are more likely to generate a group of close outlying points.
In contrast to OOS, both IOSs achieve superior performance, ranking highest among the 13 outlier detection methods. While the SUN-MCCD method showed promising results in Section 6 of [15], outperforming other clustering-based approaches, LOF achieved higher $F_{2}$-scores under most simulation settings. Fortunately, the two IOSs demonstrate substantial improvement over the SUN-MCCD method. For example, in the mixed point process, the $F_{2}$-scores of UNCCD-IOS are 0.773, 0.884, 0.917, 0.938, 0.952, 0.997, and 1.000, substantially higher than those achieved by the SUN-MCCD method (0.736, 0.816, 0.866, 0.850, 0.682, 0.535, and 0.496), particularly when $d\geq 20$. Furthermore, considering the ranking in Table 13, UNCCD-IOS outperforms LOF under most simulation settings. For example, under the Matérn cluster process, the $F_{2}$-scores of UNCCD-IOS are higher than LOF when $d\geq 5$. Consider RKCCD-IOS, although it yields worse overall performance compared to UNCCD-IOS (because UN-CCDs achieves better clustering results than RK-CCDs), it is still comparable to or better than LOF, making it the second-best method here.
@@ 4 Real Data Examples
In this section, we evaluate the performance of all four CCD-based OSs using real-life data sets, and compare them with established methods. We also want to investigate if the four OSs outperform previously CCD-based methods proposed in [15]. The data sets, obtained from Outlier Detection Data Sets (ODDS) [12] and ELKI Outlier Data Sets [14], are generally more complex than the artificial data sets used in Sections 2 and 3. Prior to outlier detection, we preprocess the data sets by normalizing all the features. Traditional normalization, subtracting the sample mean and dividing by the sample standard deviation, is not robust to outliers with extreme values [10]. To address it, we employ a robust alternative way with mean and standard deviation replaced by the median (Med) and the Normalized Median Absolute Deviation about the median (MADN), respectively.
@@
Parameter selections for the benchmark methods remains the same as in Section 3. The four OSs use a threshold of 2, and the parameter $S_{min}$ for the two IOSs is set to 0. We record TPRs, TNRs, BAs, and $F_{2}$-scores in Tables 15 and 16.
@@
Table 16: The BAs and $F_{2}$-scores of selected outlier detection methods on real-life data sets.
In the hepatitis data set, UNCCD-IOS delivers the highest $F_{2}$-score (0.541) among all the methods. The UN-MCCD and SUN-MCCD methods achieve the second-best performance with $F_{2}$-score equal to 0.446. In comparison, the MST method yields a $F_{2}$-score of 0.375. All the other methods deliver considerably lower TPRs, leading to worse performance.
@@
Consider the glass data set, the SUN-MCCD method achieves the highest performance with a $F_{2}$-score of 0.324. LOF and the MST method deliver comparable $F_{2}$-scores of 0.289 and 0.313, respectively. The four OSs, along with DBSCAN, ODIN, and iForest, can only capture a small proportion of outliers, leading to poor performance. While the U-MCCD and SU-MCCD methods successfully identify all outliers, their TNRs are only 0.363.
The two ISOs get substantially higher $F_{2}$-scores (0.515 and 0.549) than the other CCD-based methods or OOSs under the WBC data set, because both of them achieve TPR of 1 while maintaining high TNRs. However, DBSCAN and iForest perform better, with $F_{2}$-scores being 0.652 and 0.656, respectively. Other established methods perform much worse.

### v2

@@ 3.2 Under Random Cluster Process
In the previous part, we have simulated three Neyman-Scott cluster processes (the Matérn, Thomas, and mixed cluster processes) to introduce randomness in both the location and size of clusters, The four CCD-based outlier detection methods were compared to several established state-of-the-art approaches, including Local Outlier Factor (LOF) [breunig2000lof], Density Based Spatial Clustering of Applications with Noise (DBSCAN) [ester1996density], the Minimal Spanning Tree (MST) Method [MST], Outlier Detection using In-degree Number (ODIN) [hautamaki2004outlier], and isolation Forest (iForest) [liu2008isolation]. Simulation analysis shows LOF achieved the highest $F_{2}$ scores in over half of the simulation settings, and the SUN-MCCD method demonstrated the second-best overall performance while simultaneously providing clustering results. Moreover, it is worth noting that the SUN-MCCD method outperforms other cluster-based outlier detection methods by substantial margins, which makes it a decent choice under most simulation settings.
@@
We assess the performance of the four outlying scores by repeating the random cluster process in Tech Report 1. For comparison, this analysis included the four CCD-based methods and existing outlier detection methods. To expand on the previous evaluation, we increase the maximum dimensions from 20 to 100 to evaluate their performance in high-dimensional space. As noted in Section 3, IOS can be affected by the masking problem. This issue is exacerbated as the data size $n$ goes from 50 to 1000 in low-dimensional settings due to the increase in outlier intensity. To address this, we introduce a parameter, $S_{min}$, set to 0.04, for the two IOSs (excluding OOSs since they are not cluster-based), filtering the clusters with sizes below $S_{min}$. We use the same thresholds in Tables 1 and 2 for the Matérn cluster process and the Thomas cluster process, respectively. For the mixed cluster process, we set the thresholds as the mean of the values from the two tables. For example, when $d=100$, the thresholds of RKCCD-OOS are 13 (Table 1) and 10 (Table 2), resulting in a threshold of 11.5 for the mixed cluster process.
@@
For LOF, the lower and upper bounds of $k$ are 11 and 30, respectively. The highest LOF is identified for each point, with a threshold of 1.5. In DBSCAN, the outlier percentage is assumed to be known to get an appropriate cutoff value for the $4$-dist, serving as a threshold for outliers. For MST, inconsistent edges are identified using a threshold of 1.2, and clusters smaller than 2% of the dataset are labeled as outliers. As for ODIN, we set the input parameters $k$ and $T$ to $n^{0.5}$ and $n^{0.33}$, where $n$ is the data size; iForests are constructed with 1000 iTrees, each with a sub-sample size of 256, a threshold of 0.55 is applied to capture the outliers.
@@
RKCCD-OOS 0.898 0.924 0.891 0.945 0.816 0.955 0.750 0.993 0.539 0.991 0.314 0.907 0.311 0.905
RKCCD-IOS 0.956 0.973 0.952 0.977 0.976 0.966 0.991 0.984 0.997 0.990 0.997 1.000 0.999 1.000
UNCCD-OOS 0.699 0.971 0.824 0.985 0.851 0.995 0.734 0.995 0.643 0.994 0.417 0.940 0.314 0.904
UNCCD-IOS 0.949 0.953 0.955 0.970 0.958 0.979 0.985 0.978 0.988 0.984 0.998 0.999 1.000 1.000
@@
Table 7: The TPRs and TNRs of selected outlier detection methods and OSs under a Matérn cluster process (the simulation setting I in our previous work [shi2024outlier]).
@@
RKCCD-OOS 0.911 0.675 0.918 0.793 0.886 0.764 0.872 0.771 0.765 0.573 0.611 0.262 0.608 0.259
RKCCD-IOS 0.965 0.840 0.965 0.898 0.971 0.909 0.988 0.959 0.994 0.974 0.999 0.997 1.000 0.999
UNCCD-OOS 0.835 0.630 0.905 0.815 0.923 0.865 0.865 0.762 0.819 0.677 0.679 0.407 0.609 0.261
UNCCD-IOS 0.951 0.770 0.963 0.884 0.969 0.917 0.982 0.938 0.986 0.954 0.999 0.997 1.000 1.000
@@
Table 8: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under a Matérn cluster process (the simulation setting I in our previous work [shi2024outlier]).
@@
RKCCD-OOS 0.752 0.925 0.721 0.945 0.792 0.942 0.854 0.940 0.717 0.922 0.540 0.921 0.547 0.943
RKCCD-IOS 0.891 0.951 0.918 0.936 0.941 0.919 0.943 0.898 0.809 0.997 0.912 0.999 0.924 0.997
UNCCD-OOS 0.544 0.956 0.659 0.961 0.849 0.953 0.805 0.942 0.766 0.942 0.570 0.898 0.548 0.943
UNCCD-IOS 0.868 0.953 0.897 0.949 0.929 0.941 0.899 0.963 0.914 0.917 0.927 0.980 0.924 0.981
@@
Table 9: The TPRs and TNRs of selected outlier detection methods and OSs under a Thomas cluster process (the simulation setting II in our previous work [shi2024outlier]).
@@
RKCCD-OOS 0.839 0.518 0.833 0.646 0.867 0.724 0.897 0.775 0.820 0.648 0.731 0.506 0.745 0.563
RKCCD-IOS 0.921 0.678 0.927 0.782 0.930 0.804 0.921 0.783 0.903 0.831 0.956 0.925 0.961 0.931
UNCCD-OOS 0.750 0.437 0.810 0.624 0.901 0.785 0.874 0.741 0.854 0.711 0.734 0.509 0.746 0.538
UNCCD-IOS 0.911 0.667 0.923 0.789 0.935 0.826 0.931 0.844 0.916 0.786 0.954 0.905 0.953 0.915
@@
Table 10: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under Thomas cluster process (the simulation setting II in our previous work [shi2024outlier]).
@@
RKCCD-OOS 0.752 0.946 0.753 0.962 0.747 0.977 0.812 0.972 0.630 0.964 0.473 0.792 0.367 0.808
RKCCD-IOS 0.929 0.950 0.922 0.959 0.887 0.949 0.868 0.937 0.969 0.880 0.997 0.950 1.000 0.994
UNCCD-OOS 0.666 0.956 0.761 0.966 0.848 0.973 0.781 0.966 0.718 0.966 0.558 0.816 0.421 0.744
@@
Table 11: The TPRs and TNRs of selected outlier detection methods and OSs under a mixed cluster process (the simulation setting III in our previous work [shi2024outlier]).
@@
RKCCD-OOS 0.849 0.592 0.858 0.707 0.862 0.739 0.892 0.790 0.797 0.620 0.633 0.326 0.588 0.242
RKCCD-IOS 0.940 0.729 0.941 0.833 0.918 0.804 0.903 0.767 0.925 0.770 0.974 0.913 0.997 0.989
UNCCD-OOS 0.811 0.548 0.864 0.721 0.911 0.819 0.874 0.755 0.842 0.701 0.687 0.433 0.583 0.252
@@
Table 12: The BAs and $F_{2}$-scores of selected outlier detection methods and OSs under a mixed cluster process (the simulation setting III in our previous work [shi2024outlier]).
@@
U-MCCDs 8 8 9 12 11 8 8 9 9 11 12 13 7 7 9 9 10 12 13 8 7
SU-MCCDs 4 5 5 6 10 7 8 3 3 7 11 12 7 9 4 5 5 11 11 8 7
UN-MCCDs 9 7 7 5 8 11 10 7 6 8 8 11 13 13 8 8 7 5 9 10 9
@@
RKCCD-OOS 10 11 12 11 12 13 13 11 10 10 5 9 11 8 10 11 11 7 10 13 13
RKCCD-IOS 2 2 2 1 1 1 2 5 5 4 4 1 1 1 6 3 8 8 5 2 2
UNCCD-OOS 11 9 8 10 9 12 12 12 12 5 9 7 10 11 12 10 5 10 7 11 11
UNCCD-IOS 7 3 1 2 2 1 1 6 4 3 2 3 2 2 3 2 1 1 1 1 1
LOF 1 1 6 8 5 6 5 1 1 1 1 2 4 4 1 1 2 3 4 4 4
DBSCAN 3 10 10 9 7 5 6 2 8 9 7 5 5 6 2 7 9 9 6 6 6
@@
ODIN 6 4 4 4 4 4 4 8 7 6 6 6 6 5 7 6 4 4 3 5 5
iForest 12 12 11 7 3 3 3 10 11 12 10 4 3 3 11 12 12 6 2 3 3
Table 13: The rankings (by $F_{2}$-scores) of all the methods under each simulation setting of this section, top 3 are highlighted in bold. UNCCD-IOS delivers the best overall performance compared to all the other methods.
Considering the two OOSs, while achieving high TNRs, both deliver lower TPRs and $F_{2}$-scores, ranking around $10^{th}$ place out of 13 methods under most simulation settings. For instance, the TPRs of RKCCD-OOS are 0.752, 0.751, 0.792, 0.854, 0.717, 0.540, and 0.547 under the Thomas cluster process, substantially lower than most other methods. The reason stems from the limitation of OOS: OOS measures the outlyingness of a point by comparing its vicinity density to its outbound neighbors. This approach is not robust to the masking problem. Moreover, the masking problem is more severe under the simulations with the three Neyman-Scott cluster processes compared to those with distinct clusters, as the random cluster processes are more likely to generate a group of close outlying points.
In contrast to OOS, both IOSs achieve superior performance, ranking highest among the 13 outlier detection methods. While the SUN-MCCD method previously showed promising results in our previous work [shi2024outlier], outperforming other clustering-based approaches, LOF achieved higher $F_{2}$-scores under most simulation settings. Fortunately, the two IOSs demonstrate substantial improvement over the SUN-MCCD method. For example, in the mixed point process, the $F_{2}$-scores of UNCCD-IOS are 0.773, 0.884, 0.917, 0.938, 0.952, 0.997, and 1.000, substantially higher than those achieved by the SUN-MCCD method (0.736, 0.816, 0.866, 0.850, 0.682, 0.562, and 0.496), particularly when $d\geq 20$. Furthermore, considering the ranking in Table 13, UNCCD-IOS outperforms LOF under most simulation settings. For example, under the Matérn cluster process, the $F_{2}$-scores of UNCCD-IOS are higher than LOF when $d\geq 5$. Consider RKCCD-IOS, although it yields worse overall performance compared to UNCCD-IOS (because UN-CCDs achieve better clustering results than RK-CCDs), it is still comparable to or better than LOF, making it the second-best method here.
@@ 4 Real Data Examples
In this section, we evaluate the performance of all four CCD-based OSs using real-life datasets, and compare them with established methods. We also want to investigate if the four OSs outperform previously CCD-based methods proposed in Tech Report 1. The datasets, obtained from Outlier Detection Datasets (ODDS) [Rayana2016] and ELKI Outlier Datasets [DBLP:journals/corr/abs-1902-03616], are generally more complex than the artificial datasets used in Sections 3. Prior to outlier detection, we preprocess the datasets by normalizing all the features. Traditional normalization, subtracting the sample mean and dividing by the sample standard deviation, is not robust to outliers with extreme values [maronna2019robust]. To address it, we employ a robust alternative way with mean and standard deviation replaced by the median (Med) and the Normalized Median Absolute Deviation about the median (MADN), respectively.
@@
Parameter selections for the benchmark methods remain the same as in Section 3.2. The four OSs use a threshold of 2, and the parameter $S_{min}$ for the two IOSs is set to 0. We record TPRs, TNRs, BAs, and $F_{2}$-scores in Tables 15 and 16.
@@
Table 16: The BAs and $F_{2}$-scores of selected outlier detection methods on real-life datasets. In general, the two IOS not only outperform previous CCD-based methods but also surpass established methods.
In the hepatitis dataset, UNCCD-IOS delivers the highest $F_{2}$-score (0.541) among all the methods. The UN-MCCD and SUN-MCCD methods achieve the second-best performance with $F_{2}$-score equal to 0.446. In comparison, the MST method yields an $F_{2}$-score of 0.375. All the other methods deliver considerably lower TPRs, leading to worse performance.
@@
Consider the glass dataset, the SUN-MCCD method achieves the highest performance with an $F_{2}$-score of 0.324. LOF and the MST method deliver comparable $F_{2}$-scores of 0.289 and 0.313, respectively. The four OSs, along with DBSCAN, ODIN, and iForest, can only capture a small proportion of outliers, leading to poor performance. While the U-MCCD and SU-MCCD methods successfully identify all outliers, their TNRs are only 0.363.
The two ISOs get substantially higher $F_{2}$-scores (0.515 and 0.549) than the other CCD-based methods or OOSs under the WBC dataset, because both of them achieve a TPR of 1 while maintaining high TNRs. However, DBSCAN and iForest perform better, with $F_{2}$-scores being 0.652 and 0.656, respectively. Other established methods perform much worse.